# Minimax Hypothesis Testing for the Bradley-Terry-Luce Model

### v1

@@ I-C Outline
I-D Notational Preliminaries
We briefly collect some notation here that is used throughout this work. We let ${\mathbb{R}}=(-\infty,\infty),{\mathbb{R}}_{+}=(0,\infty)$ and ${\mathbb{N}}\triangleq\{1,2,3,\dots\}$ denote the sets of real and natural numbers, respectively. For any $n\in{\mathbb{N}}$, we let ${\mathbb{R}}^{n}$ and ${\mathbb{R}}^{n\times n}$ be the sets of all $n$-length vectors and $n\times n$ matrices, respectively, ${\mathbf{1}}_{n}\in{\mathbb{R}}^{n}$ be the column vector with all entries equal to $1$, and $[n]\triangleq\{1,\dots,n\}$. For any (index) set ${\mathcal{S}}\subseteq[n]$, we let ${\mathbf{1}}_{\mathcal{S}}\in{\mathbb{R}}^{n}$ denote a vector in which the entries are $1$ for elements in the set ${\mathcal{S}}$ and 00 otherwise. Moreover, we let ${\mathcal{P}}_{n}$ denote the $(n-1)$-dimensional probability simplex in ${\mathbb{R}}^{n}$, and ${\mathbbm{1}}_{\mathcal{A}}$ denote the indicator function on the set $\mathcal{A}$. Next, for any vector $x\in{\mathbb{R}}^{n}$, $\|x\|_{2}$ and $\|x\|_{\infty}$ denote its $\ell^{2}$- and $\ell^{\infty}$-norms, $x^{p}$ denotes the entrywise $p$th power of $x$, i.e., $x^{p}=[x_{1}^{p},x_{2}^{p},\cdots,x_{n}^{p}]^{\mathrm{T}}$, and $\operatorname{{diag}}(x)\in{\mathbb{R}}^{n\times n}$ is the diagonal matrix with $x$ along its principal diagonal. For any matrix $A\in{\mathbb{R}}^{n\times n}$, $\|A\|_{2}$, $\|A\|_{\mathrm{F}}$, and $\text{Tr}(A)$ denote the standard spectral norm, Frobenius norm, and trace of $A$, respectively, $A_{ij}$ denotes the $(i,j)$th element of $A$, $A_{:,i}$ denotes the $i$th column of $A$, and $A_{j,:}$ denotes the transpose of the $j$th row of $A$ (i.e., in column vector form). Let $\sigma_{1}(A)\geq\sigma_{2}(A)\geq\dots\geq\sigma_{n}(A)$ represent the (ordered) singular values of $A\in{\mathbb{R}}^{n\times n}$. For an entrywise strictly positive vector $\pi\in{\mathbb{R}}^{n}$, we define a Hilbert space $\ell^{2}(\pi)$ on ${\mathbb{R}}^{n}$ with inner product $\langle x,y\rangle_{\pi}=\sum_{i=1}^{n}\pi_{i}x_{i}y_{i}$ and corresponding vector, operator, and weighted Frobenius norms $\|x\|_{\pi}=\sqrt{\langle x,x\rangle_{\pi}}$, $\|A\|_{\pi}=\sup_{\|x\|_{\pi}=1}\|x^{\mathrm{T}}A\|_{\pi}$, and $\|A\|_{\pi,\mathrm{F}}=\big{(}\sum_{i}\sum_{j}\pi_{j}A^{2}_{ij}\big{)}^{1/2}$, respectively. Finally, we utilize standard Bachmann-Landau asymptotic notation, e.g., $O(\cdot)$, where it is assumed that the parameter $n\rightarrow\infty$ or the parameter $k\rightarrow\infty$, where $n$ represents the number of items and $k$ is the number of pairwise comparisons per pair of items. Throughout this paper, we use “high probability” to refer to a probability of at least $1-1/\text{poly}(n)$, where $\text{poly}(n)$ captures a polynomial function of $n$.
II Formal Setup and Goal
We begin by introducing a general pairwise comparison model. Consider a scenario where a set of $n$ agents, indexed by $[n]$ with $\1n\in{\mathbb{N}}\backslash\!\{1\}$, engage in a tournament consisting of several observed pairwise comparisons. This scenario is ubiquitous in many application domains, such as sports tournaments, discrete choice models in economics, etc. For example, in a sports tournament, $[n]$ represents the teams or players that play pairwise games with each other. Similarly, in the discrete choice models in economics, $[n]$ represents the available alternatives that an individual may choose from.
@@ II Formal Setup and Goal
Within this context, we assume that each observation corresponds to a pairwise comparison between agents $i$ and $j$ for distinct $i,j\in[n].$ We focus on a general asymmetric setting, where an “$i$ vs. $j$” comparison can have a different interpretation to a “$j$ vs. $i$” comparison. For example, in an “$i$ vs. $j$” comparison, $i$ and $j$ may represent the home-team and the away-team, respectively. This type of asymmetry is common in sports like cricket, soccer, etc. [22], where it is often observed that the home-team has a certain advantage over the away-team. Nevertheless, it is worth mentioning that all our ensuing theorems, insights, and other results can be easily extended to the symmetric setting, where “$i$ vs. $j$” and “$j$ vs. $i$” comparisons are considered equivalent.
@@
In general, comparisons between all pairs $i,j\in[n]$ may not be observed. To model this, we assume that we are given an induced observation graph ${\mathcal{G}}=([n],{\mathcal{E}})$ with vertex set $[n]$ and edge set ${\mathcal{E}}$, where an edge $(i,j)\in{\mathcal{E}}$ (with $i\neq j$) exists if and only if comparisons of the form “$i$ vs. $j$” are observed. Additionally, for convenience, self-loops are included in ${\mathcal{E}}$ for all vertices (i.e., $(i,i)\in{\mathcal{E}}$ for all $i\in[n]$). Let $E\in\{0,1\}^{n\times n}$ be the adjacency matrix of the graph ${\mathcal{G}}$ with $E_{ij}=1$ if $(i,j)\in{\mathcal{E}}$ and 00 otherwise. We also define the projection operator ${\mathcal{P}_{\mathcal{E}}}(X)\triangleq X\odot E$ for $X\in{\mathbb{R}}^{n\times n}$, where $\odot$ denotes the Hadamard product. Moreover, we assume that the edge set ${\mathcal{E}}$ is symmetric (i.e., the graph is actually undirected); hence, if “$i$ vs. $j$” comparisons are observed, then we require that “$j$ vs. $i$” comparisons are observed as well. Finally, we assume that the graph is connected and is given a priori, i.e., it does not depend on the outcomes of observed pairwise comparisons. (Connectedness of the graph is a required, standard assumption in the literature, cf. [55, 7].) In the sequel, we will further specify several classes of graphs where we can prove theoretical guarantees on hypothesis testing for the BTL model.
@@
The next definition presents the general pairwise comparison model. Note that in the literature, various well-known probabilistic models have been developed to capture pairwise comparison settings, such as the BTL model [2, 3, 5], Thurstonian model [24], non-parametric models [27, 28], etc. These are all specializations of the general model below.
@@
For any pair of distinct agents $i,j\in[n]$, let $p_{ij}\in(0,1)$ denote the probability that agent $j$ beats agent $i$ in a “$i$ vs. $j$” pairwise comparison. We refer to the collection of parameters $\{p_{ij}:(i,j)\in{\mathcal{E}},\,i\neq j\}$ as a pairwise comparison model (on ${\mathcal{E}}$).
@@
where we have set $P_{ii}=\frac{1}{2}$ for notational convenience. In our analysis, we will find it convenient to assign a time-homogenous Markov chain (or a row stochastic matrix) on the finite state space $[n]$ to any pairwise comparison model. We establish this canonical assignment as follows.
@@
$S_{ij}\triangleq\begin{cases}\displaystyle{\frac{p_{ij}}{d}},&i\neq j{\text{ and }}(i,j)\in{\mathcal{E}}\\ \displaystyle{1-\frac{1}{d}\sum_{\begin{subarray}{c}k\in[n]\backslash\!\{i\}:\\ (i,k)\in{\mathcal{E}}\end{subarray}}{p_{ik}}},&i=j\\ 0,&\text{otherwise}\end{cases},$ (3)
@@
For fixed induced graphs, we set the parameter $d=2d_{\max},$ where $d_{\max}\triangleq\max_{i\in[n]}\sum_{j=1}^{n}E_{ij}$ is the maximum degree of a node in ${\mathcal{G}}$. Further discussion regarding this parameter can be found in Section III-E. As noted earlier, the most well-known specialization of the pairwise comparison model in Definition 1 is the BTL model defined below [2, 3, 5].
Definition 3 (BTL Model [2, 3, 5]).
@@
$\forall(i,j)\in{\mathcal{E}}\textup{ with }i\neq j,\,\enspace p_{ij}=\frac{\alpha_{j}}{\alpha_{i}+\alpha_{j}}.$
@@
Next, we explain the data generation process. Fix any pairwise comparison model $\{p_{ij}\in(0,1):(i,j)\in{\mathcal{E}},\,i\neq j\}$.
For any pair $i\neq j$, define the outcome of the $m$th “$i$ vs. $j$” pairwise comparison between them as a Bernoulli random variable
$Z_{m,{ij}}\triangleq\begin{cases}1,&\text{ if }j\text{ beats }i\ (\text{with probability }p_{ij})\\ 0,&\text{ if }i\text{ beats }j\ (\text{with probability }1-p_{ij})\end{cases},$ (4)
for $m\in[k_{ij}]$, where $k_{ij}$ denotes the number of observed “$i$ vs. $j$” comparisons. We assume throughout that the observation random variables ${\mathcal{Z}}\triangleq\{Z_{m,{ij}}:(i,j)\in{\mathcal{E}},\,i\neq j,\,m\in[k_{i,j}]\}$ are mutually independent. Moreover, let $Z_{ij}\triangleq\sum_{m=1}^{k_{ij}}Z_{m,{ij}}$. Clearly, it follows that for any $(i,j)\in{\mathcal{E}}$ with $i\neq j$, $Z_{ij}$ is a binomial random variable, i.e., $Z_{ij}\sim\text{Bin}(k_{ij},p_{ij})$, and for simplicity, we set $Z_{ii}=k_{ii}=0$ for all $i\in[n]$.
@@ II-A Main Goal
[…]
@@
Additionally, for analytical tractability, we make the following standard assumption on the pairwise comparison models under consideration (see [10, 11, 15]).
Assumption 1 (Dynamic Range).
We assume that there is a constant $\delta\in(0,1)$ such that for all $(i,j)\in{\mathcal{E}}$,
$\frac{\delta}{1+\delta}\leq p_{ij}\leq\frac{1}{1+\delta}.$ (6)
To pose the hypothesis testing problem in Eq. 5 more rigorously, we demonstrate an interesting relation between a BTL model and its canonical Markov matrix. Recall that a Markov chain on the state space $[n]$, defined by the row stochastic matrix $W\in{\mathbb{R}}^{n\times n}$, is said to be reversible if it satisfies the detailed balance conditions [56, Section 1.6]:
$\forall i,j\in[n],\,i\neq j,\enspace\pi_{i}W_{ij}=\pi_{j}W_{ji},$ (7)
@@
It turns out that there is a tight connection between reversible Markov chains and the BTL model. This is elucidated in the ensuing proposition, which is a more general version of [57, Lemma 6], [10, Section 2.2].
@@
II-B Expansion Properties of ${\mathcal{G}}$
In addition to Assumption 1, we need a few more assumptions on the underlying graph ${\mathcal{G}}$ and the probabilities $p_{ij}$ of the pairwise comparison model. Broadly, these assumptions necessitate that the graph ${\mathcal{G}}$ has edge expansion properties and is almost regular. To present the details, we start by considering the divergence transition matrix (DTM) [58, 59, 60],
$R=\Pi^{1/2}S\Pi^{-1/2}\,,$ (9)
corresponding to a canonical Markov matrix $S$ associated with a valid pairwise comparison matrix $P$, where $\pi$ is the stationary distribution of $S$ and $\Pi\triangleq\operatorname{{diag}}(\pi)$. We also note that due to Assumption 1, the connectedness of ${\mathcal{G}}$, and the aperiodicity of $S$ (since $S_{ii}>0$ for all $i\in[n]$), $S$ possesses a unique, entrywise strictly positive, stationary distribution $\pi$. This ensures that the quantities we introduce in the sequel, such as edge expansion and principal ratio, are well-defined.
Instead of directly imposing assumptions on the expansion properties of ${\mathcal{G}}$, we impose an assumption on the edge expansion of the DTM $R$ for simplicity. Observing that $R$ possesses a Perron-Frobenius eigenvalue [61, 62] of $1$ with corresponding left and right eigenvectors of $\sqrt{\pi}$ that are entrywise strictly positive [58, Proposition 2.2], we define the edge expansion of $R$ as follows.
Definition 4 (Edge Expansion of Non-negative Matrices [63]).
@@ II-B Expansion Properties of ${\mathcal{G}}$
Consider any non-negative matrix $R\in\mathbb{R}^{n\times n}$ with a Perron-Frobenius eigenvalue of $1$ and corresponding left and right eigenvectors $u$ and $v$, which we assume are entrywise strictly positive and normalized such that $u^{\mathrm{T}}v=1$. Let $D_{u}\triangleq\operatorname{{diag}}(u)$ and $D_{v}\triangleq\operatorname{{diag}}(v)$ be diagonal matrices with $u$ and $v$ on the principal diagonals, respectively. Then, the edge expansion of $R$ is defined as
$\phi(R)\triangleq\min_{{\mathcal{S}}\subseteq[n]}\frac{\mathbf{1}_{{\mathcal{S}}}^{\mathrm{T}}D_{u}RD_{v}\mathbf{1}_{{\mathcal{S}}^{\mathrm{c}}}}{\min\{\mathbf{1}_{{\mathcal{S}}}^{\mathrm{T}}D_{u}D_{v}\mathbf{1}_{n},\mathbf{1}_{{\mathcal{S}}^{\mathrm{c}}}^{\mathrm{T}}D_{u}D_{v}\mathbf{1}_{n}\}}=\min_{{\mathcal{S}}\subseteq[n]}\frac{\sum_{i\in{\mathcal{S}},j\in{\mathcal{S}}^{\mathrm{c}}}R_{ij}u_{i}v_{j}}{\min\{\sum_{i\in{\mathcal{S}}}u_{i}v_{i},\sum_{i\in{\mathcal{S}}^{\mathrm{c}}}u_{i}v_{i}\}}.$ (10)
@@
Note that in cases where there are no strictly positive left and right eigenvectors $u$ and $v$ corresponding to the eigenvalue $1$, the edge expansion $\phi(R)$ is defined to be 00. For the DTM $R=\Pi^{1/2}S\Pi^{-1/2},$ its edge expansion simplifies to
$\phi(R)=\min_{{\mathcal{S}}\subseteq[n]}\frac{\sum_{i\in{\mathcal{S}},j\in{\mathcal{S}}^{\mathrm{c}}}p_{ij}\pi_{i}\mathbbm{1}_{(i,j)\in{\mathcal{E}}}}{d\min\{\sum_{i\in{\mathcal{S}}}\pi_{i},\sum_{i\in{\mathcal{S}}^{\mathrm{c}}}\pi_{i}\}}.$ (11)
We introduce the following assumption on the edge expansion of the DTM $R$ in Eq. 11.
Assumption 2 (Large Edge Expansion of DTM).
@@
For the DTM $R=\Pi^{1/2}S\Pi^{-1/2}$ corresponding to the canonical Markov matrix $S$, we assume that there exists a constant $\xi>0$ such that
$\phi(R)\geq\xi.$ (12)
As mentioned earlier, in addition to edge expansion, we require the canonical Markov matrix $S$ to be almost regular. The notion of regularity is quantified through two criteria: regularity of the underlying graph ${\mathcal{G}}$ and the principal ratio of $S$ (defined below). The regularity of the graph ${\mathcal{G}}$ is formalized in the following assumption.
Assumption 3 (Almost Regular Graph).
Let $d_{\min}\triangleq\min_{i\in[n]}\sum_{j\in[n]\backslash\!\{i\}}E_{ij}$ and $d_{\max}\triangleq\max_{i\in[n]}\sum_{j\in[n]}E_{ij}.$ We require that the graph ${\mathcal{G}}$ is connected and almost regular in the sense that there exists a constant $\kappa>0$ such that
$\frac{d_{\max}}{d_{\min}}\leq\kappa.$ (13)
Intuitively, Assumption 3 states that $d_{\max}\approx d_{\min}$, i.e., the minimum and maximum degrees are close to each other. Another measure of regularity for a non-negative matrix is the well-known concept of principal ratio, cf. [64] (which is the Birkhoff norm of the Perron-Frobenius vector [65], and is sometimes called the height of the Perron-Frobenius vector [66]).
Definition 5 (Principal Ratio).
@@
We note that the principal ratio is both a function of the underlying graph ${\mathcal{G}}$ and the pairwise comparison probabilities $p_{ij}$. Intuitively, a smaller value of this ratio, i.e., closer to $1$, signifies a “more regular” matrix $P$ (see, e.g., [64]). To proceed, we make the following assumption on the principal ratio.
@@
There exists a constant $h>0$ such that the principal ratio $h_{\pi}$ is bounded: [^1] The constants $\xi$, $\kappa$, and $h$ in Assumptions 2, 3 and 4 do not depend on $n$.
@@
It is worth mentioning that coupled with Assumptions 1 and 4, Assumption 2 can be imposed by lower bounding the canonical notion of edge expansion of ${\mathcal{G}}$. Indeed, observe that
$\phi(R)=\min_{{\mathcal{S}}\subseteq[n]}\frac{\sum_{i\in{\mathcal{S}},j\in{\mathcal{S}}^{\mathrm{c}}}p_{ij}\pi_{i}\mathbbm{1}_{(i,j)\in{\mathcal{E}}}}{d\min\{\sum_{i\in{\mathcal{S}}}\pi_{i},\sum_{i\in{\mathcal{S}}^{\mathrm{c}}}\pi_{i}\}}\geq\frac{\delta}{dh_{\pi}(1+\delta)}\min_{{\mathcal{S}}\subseteq[n]}\frac{|{\mathcal{E}}({\mathcal{S}},{\mathcal{S}}^{\mathrm{c}})|}{\min\{|{\mathcal{S}}|,|{\mathcal{S}}^{\mathrm{c}}|\}}\geq\frac{\delta}{dh(1+\delta)}\tilde{\phi}({\mathcal{G}}),$ (16)
where ${\mathcal{E}}({\mathcal{S}},{\mathcal{S}}^{\mathrm{c}})\triangleq\{(i,j)\in{\mathcal{E}}|i\in{\mathcal{S}},j\in{\mathcal{S}}^{\mathrm{c}}\}$ is the set of edges in ${\mathcal{E}}$ from set ${\mathcal{S}}$ to ${\mathcal{S}}^{\mathrm{c}}$ and $\tilde{\phi}({\mathcal{G}})$ is the usual definition of edge expansion of ${\mathcal{G}}$:
@@
automatically satisfy both Assumptions 1 and 4 (with $h=\delta^{-1}$). [^2] We refer readers to Section IV-A for an explicit expression for the stationary distribution in BTL models. Furthermore, if the underlying graph ${\mathcal{G}}$ exhibits significant edge expansion in the canonical sense, i.e., $\tilde{\phi}({\mathcal{G}})\geq\tilde{\epsilon}d$ for some constant $\tilde{\epsilon}>0$, and adheres to Assumption 3, then all four assumptions are simultaneously satisfied for all BTL models with bounded condition number. For general pairwise comparison models, we will present several classes of graphs in Section III-E, e.g., complete graphs, dense regular graphs, and Erdős-Rényi random graphs, where any pairwise comparison model satisfying Assumption 1 will also satisfy Assumptions 2, 3 and 4 (with high probability, where appropriate).

### v2

@@ II Formal Setup and Goal
Within this context, we assume that each observation corresponds to a pairwise comparison between agents $i$ and $j$ for distinct $i,j\in[n]$. We focus on a general asymmetric setting, where an “$i$ vs. jj” comparison can have a different interpretation to a “$j$ vs. ii” comparison. For example, in an “$i$ vs. jj” comparison, $i$ and $j$ may represent the home team and the away team, respectively. This type of asymmetry is common in sports like cricket, soccer, etc. [Morley2005], where it is often observed that the home team has a certain advantage over the away team. Nevertheless, it is worth mentioning that many of our ensuing theorems, insights, and other results can be extended with some effort to the symmetric setting, where “$i$ vs. jj” and “$j$ vs. ii” comparisons are considered equivalent.
@@
In general, comparisons between all pairs $i,j\in[n]$ may not be observed. To model this, we assume that we are given an induced observation graph ${\mathcal{G}}=([n],{\mathcal{E}})$ with vertex set $[n]$ and edge set $\mathcal{E}$}, where an edge $(i,j)\in{\mathcal{E}}$ (with $i\neq j$) exists if and only if comparisons of the form “$i$ vs. jj” are observed. Additionally, for convenience, self-loops are included in $\mathcal{E}$} for all vertices (i.e., $(i,i)\in{\mathcal{E}}$ for all $i\in[n]$). Let $E\in\{0,1\}^{n\times n}$ be the adjacency matrix of the graph $\mathcal{G}$} with $E_{ij}=1$ if $(i,j)\in{\mathcal{E}}$ and 0 otherwise. We also define the projection operator ${\mathcal{P}_{\mathcal{E}}}(X)\triangleq X\odot E$ for $X\in{\mathbb{R}}^{n\times n}$, where $\odot$ denotes the Hadamard product. Moreover, we assume that the edge set $\mathcal{E}$} is symmetric (i.e., the graph is actually undirected); hence, if “$i$ vs. jj” comparisons are observed, then we require that “$j$ vs. ii” comparisons are observed as well. Finally, we assume that the graph is connected and is given a priori, i.e., it does not depend on the outcomes of observed pairwise comparisons. (Connectedness of the graph is a required, standard assumption in the literature, cf. [Ford1957, Hunter2004].) In the sequel, we will further specify several classes of graphs where we can prove theoretical guarantees on hypothesis testing for the BTL model.
@@
The next definition presents the general pairwise comparison model. Note that in the literature, various well-known probabilistic models have been developed to capture pairwise comparison settings, such as the BTL model [BradleyTerry1952, Luce1959, McFadden1973], Thurstonian model [Thurstone1927], non-parametric models [Chatterjee2015, Shahetal2016], etc. These are all specializations of the general model below.
@@
For any pair of distinct agents $i,j\in[n]$, let $p_{ij}\in(0,1)$ denote the probability that agent $j$ beats agent $i$ in a “$i$ vs. jj” pairwise comparison. We refer to the collection of parameters $\{p_{ij}:(i,j)\in{\mathcal{E}},\,i\neq j\}$ as a pairwise comparison model (on $\mathcal{E}$}).
@@
where we have set $P_{ii}=\frac{1}{2}$ for notational convenience. In our analysis, we find it convenient to assign a time-homogeneous Markov chain (or a row stochastic matrix) on the finite state space $[n]$ to any pairwise comparison model. We establish this canonical assignment as follows.
@@
$S_{ij}\triangleq\begin{cases}\displaystyle{\frac{p_{ij}}{d}},&i\neq j{\text{ and }}(i,j)\in{\mathcal{E}}\\ \displaystyle{1-\frac{1}{d}\sum_{\begin{subarray}{c}u\in[n]\backslash\!\{i\}:\\ (i,u)\in{\mathcal{E}}\end{subarray}}{p_{iu}}},&i=j\\ 0,&\text{otherwise}\end{cases},$ (3)
@@
For fixed induced graphs, we set the parameter $d=2d_{\max},$ where $d_{\max}\triangleq\max_{i\in[n]}\sum_{j=1}^{n}E_{ij}$ is the maximum degree of a node in $\mathcal{G}$}. Further discussion regarding this parameter can be found in Section III-E. We note that due to the strict positivity of $p_{ij}$’s and connectedness of $\mathcal{G}$}, $S$ possesses a unique, entrywise strictly positive, stationary distribution. As noted earlier, the most well-known specialization of the pairwise comparison model in Definition 1 is the BTL model defined below [BradleyTerry1952, Luce1959, McFadden1973].
Definition 3 (BTL Model [BradleyTerry1952, Luce1959, McFadden1973]).
@@
$\forall(i,j)\in{\mathcal{E}}\textup{ with }i\neq j,\,\kern 5.0ptp_{ij}=\frac{\alpha_{j}}{\alpha_{i}+\alpha_{j}}.$
@@
Next, we explain the data generation process. Fix any pairwise comparison model $\{p_{ij}\in(0,1):(i,j)\in{\mathcal{E}},\,i\neq j\}$. For any pair $i\neq j$, define the outcome of the $m$th “$i$ vs. jj” pairwise comparison between them as a Bernoulli random variable
[…]
for $m\in[k_{ij}]$, where $k_{ij}$ denotes the number of observed “$i$ vs. jj” comparisons. We assume throughout that the observation random variables ${\mathcal{Z}}\triangleq\{Z_{m,{ij}}:(i,j)\in{\mathcal{E}},\,i\neq j,\,m\in[k_{ij}]\}$ are mutually independent. Moreover, let $Z_{ij}\triangleq\sum_{m=1}^{k_{ij}}Z_{m,{ij}}$. Clearly, it follows that for any $(i,j)\in{\mathcal{E}}$ with $i\neq j$, $Z_{ij}$ is a binomial random variable, i.e., $Z_{ij}\sim\text{Bin}(k_{ij},p_{ij})$, and for simplicity, we set $Z_{ii}=k_{ii}=0$ for all $i\in[n]$.
@@ II-A Main Goal
$H_{1}$ $:{\mathcal{Z}}\sim\text{pairwise comparison model that is not BTL},$
@@
To pose the hypothesis testing problem in Eq. 5 more rigorously, we demonstrate an interesting relation between a BTL model and its canonical Markov matrix. Recall that a Markov chain on the state space $[n]$, defined by the row stochastic matrix $W\in{\mathbb{R}}^{n\times n}$, is said to be reversible if it satisfies the detailed balance conditions [LevinPeresWilmer2009, Section 1.6]:
$\forall i,j\in[n],\,i\neq j,\kern 5.0pt\pi_{i}W_{ij}=\pi_{j}W_{ji},$ (6)
@@
It turns out that there is a tight connection between reversible Markov chains and the BTL model. This is elucidated in the ensuing proposition, which is a slightly more general version of [RajkumarAgarwal2014, Lemma 6], [NegahbanOhShah2017, Section 2.2].
@@
II-B Expansion and Principal Ratio
To carry out our analysis, we need a few more assumptions on the underlying graph $\mathcal{G}$} and the probabilities $p_{ij}$ of the pairwise comparison model. Broadly, these assumptions ensure that the canonical Markov matrix $S$ possesses sufficient connectivity, quantified by edge expansion, and a bounded principal ratio. We first recall the notion of edge expansion of non-negative matrices, which characterizes the extremal “connectivity” or “flow” from one subset of vertices to another under stationarity.
Definition 4 (Edge Expansion of Non-negative Matrices [MehtaSchulman]).
@@ II-B Expansion and Principal Ratio
Consider any non-negative matrix $R\in\mathbb{R}^{n\times n}$ with a Perron-Frobenius eigenvalue of 11 and corresponding left and right eigenvectors $u$ and $v$, which we assume are entrywise strictly positive and normalized such that $u^{\mathrm{T}}v=1$. Let $D_{u}\triangleq\operatorname{{diag}}(u)$ and $D_{v}\triangleq\operatorname{{diag}}(v)$ be diagonal matrices with $u$ and $v$ on the principal diagonals, respectively. Then, the edge expansion of $R$ is defined as
$\phi(R)$ $\triangleq\min_{{\mathcal{S}}\subseteq[n]}\frac{\mathbf{1}_{{\mathcal{S}}}^{\mathrm{T}}D_{u}RD_{v}\mathbf{1}_{{\mathcal{S}}^{\mathrm{c}}}}{\min\{\mathbf{1}_{{\mathcal{S}}}^{\mathrm{T}}D_{u}D_{v}\mathbf{1}_{n},\mathbf{1}_{{\mathcal{S}}^{\mathrm{c}}}^{\mathrm{T}}D_{u}D_{v}\mathbf{1}_{n}\}}$
$=\min_{{\mathcal{S}}\subseteq[n]}\frac{\sum_{i\in{\mathcal{S}},j\in{\mathcal{S}}^{\mathrm{c}}}R_{ij}u_{i}v_{j}}{\min\{\sum_{i\in{\mathcal{S}}}u_{i}v_{i},\sum_{i\in{\mathcal{S}}^{\mathrm{c}}}u_{i}v_{i}\}}.$ (8)
@@
Note that in cases where there are no strictly positive left and right eigenvectors $u$ and $v$ corresponding to the eigenvalue $1$, the edge expansion $\phi(R)$ is defined to be 0. For the row-stochastic matrix $S$ with stationary distribution $\pi$, its edge expansion simplifies to
$\phi(S)=\min_{{\mathcal{S}}\subseteq[n]}\frac{\sum_{i\in{\mathcal{S}},j\in{\mathcal{S}}^{\mathrm{c}}}p_{ij}\pi_{i}\mathbbm{1}_{(i,j)\in{\mathcal{E}}}}{d\min\{\sum_{i\in{\mathcal{S}}}\pi_{i},\sum_{i\in{\mathcal{S}}^{\mathrm{c}}}\pi_{i}\}}.$ (9)
We introduce the following assumption on the edge expansion of $S$.
Assumption 1 (Edge Expansion).
@@
For the canonical Markov matrix $S$ corresponding to the pairwise comparison model $P$, we assume that there exists a constant $\xi>0$ such that
$\phi(S)\geq\xi.$ (10)
As mentioned earlier, in addition to edge expansion, we require the canonical Markov matrix $S$ to have a bound on its principal ratio, cf. [EigenvectorsIrregularGraphs] (which is the Birkhoff norm of the Perron-Frobenius vector [BirkhoffContractionCoefficient], and is sometimes called the height of the Perron-Frobenius vector [Minc1970]).
Definition 5 (Principal Ratio [EigenvectorsIrregularGraphs]).
@@
We note that the principal ratio is both a function of the underlying graph $\mathcal{G}$} and the pairwise comparison probabilities $p_{ij}$. Intuitively, it captures a notion of “regularity” of Markov matrices like $S$ (akin to the regularity of vanilla graphs like $\mathcal{G}$}). Indeed, a principal ratio closer to $1$ signifies a “more regular” Markov matrix $S$; for instance, in random walks on graphs, the principal ratio equals $1$ if and only if the graph is regular (i.e., all nodes have identical degrees) [EigenvectorsIrregularGraphs]. This ratio has a long history in matrix theory and the analysis of Markov chains [birkhoff1, birkhoff2]. For example, Birkhoff’s geometric ergodicity theory connects the principal ratio to the contraction coefficient in projective metrics, where smaller ratios imply faster convergence to stationarity [BirkhoffContractionCoefficient]. Since the principal ratio impacts the rate of convergence of random walks on graphs, it also arises in the analysis of spectral algorithms, such as PageRank [PagerankPrincipalRatio], and matrix analysis [ostrowski1952]. Furthermore, several combinatorial studies have derived bounds on the principal ratio under various settings [EigenvectorsIrregularGraphs, AksoyChung2016].
To proceed, we make the following assumption on the principal ratio of $S$.
@@
There exists a constant $h>0$ such that the principal ratio $h_{\pi}$ is bounded as:
@@
We remark that we only require Assumptions 1 and 2 in the analysis of our hypothesis testing framework. [^1] We generally think of $\xi$ and $h$ in Assumptions 1 and 2 as constants that do not depend on $n$. However, our results illustrate precise dependencies with respect to these constants in cases where they depend on other problem parameters.
In the sequel, we introduce a more palatable simplifying assumption that will be further discussed. This assumption is standard in the analysis of pairwise comparison models under consideration (see, e.g., [NegahbanOhShah2017, Chenetal2019, JadbabaieMakurShah2020journal]).
Assumption 3 (Dynamic Range).
We assume that there exists a constant $\delta\in(0,1)$ such that for all $(i,j)\in{\mathcal{E}}$,
$\frac{\delta}{1+\delta}\leq p_{ij}\leq\frac{1}{1+\delta}.$ (13)
We will highlight that Assumption 3 implies Assumptions 1 and 2 for various well-known classes of induced graphs. For now, note that coupled with Assumptions 3 and 2, Assumption 1 can be expressed in terms of the canonical notion of edge expansion of $\mathcal{G}$}. Indeed, observe that
$\phi(S)$ $=\min_{{\mathcal{S}}\subseteq[n]}\frac{\sum_{i\in{\mathcal{S}},j\in{\mathcal{S}}^{\mathrm{c}}}p_{ij}\pi_{i}\mathbbm{1}_{(i,j)\in{\mathcal{E}}}}{d\min\{\sum_{i\in{\mathcal{S}}}\pi_{i},\sum_{i\in{\mathcal{S}}^{\mathrm{c}}}\pi_{i}\}}$
$\geq\frac{\delta}{dh_{\pi}(1+\delta)}\min_{{\mathcal{S}}\subseteq[n]}\frac{|{\mathcal{E}}({\mathcal{S}},{\mathcal{S}}^{\mathrm{c}})|}{\min\{|{\mathcal{S}}|,|{\mathcal{S}}^{\mathrm{c}}|\}}$
$\geq\frac{\delta}{dh(1+\delta)}\tilde{\phi}({\mathcal{G}}),$ (14)
where ${\mathcal{E}}({\mathcal{S}},{\mathcal{S}}^{\mathrm{c}})\triangleq\{(i,j)\in{\mathcal{E}}:i\in{\mathcal{S}},\,j\in{\mathcal{S}}^{\mathrm{c}}\}$ is the set of edges in $\mathcal{E}$} from set $\mathcal{S}$} to ${\mathcal{S}}^{\mathrm{c}}$, and $\tilde{\phi}({\mathcal{G}})$ is the usual definition of edge expansion of $\mathcal{G}$} [AlonChung1989]:
@@
automatically satisfy both Assumptions 3 and 2 (with $h=\delta^{-1}$). [^2] We refer readers to Section IV-A for an explicit expression for the stationary distribution in BTL models. Furthermore, if the underlying graph $\mathcal{G}$} exhibits significant edge expansion in the canonical sense, i.e., $\tilde{\phi}({\mathcal{G}})\geq\tilde{\epsilon}d$ for some constant $\tilde{\epsilon}>0$, then Assumption 1 is satisfied with $\xi=O(1)$ for all BTL models with bounded condition number. For general pairwise comparison models, we will present several classes of graphs in Section III-E, e.g., complete graphs, dense regular graphs, and Erdős-Rényi random graphs, where any pairwise comparison model satisfying Assumption 3 will also satisfy Assumptions 1 and 2 with constant $h$ and $\xi$ (with high probability, where appropriate).